### v2

# Photovoltaic power forecasting using quantum machine learning

Abstract
Predicting solar panel power output is crucial for advancing the transition to renewable energy but is complicated by the variable and non-linear nature of solar energy. This is influenced by numerous meteorological factors, geographical positioning, and photovoltaic cell properties, posing significant challenges to forecasting accuracy and grid stability. Our study introduces a suite of solutions centered around hybrid quantum neural networks designed to tackle these complexities. The first proposed model, the Hybrid Quantum Long Short-Term Memory, surpasses all tested models by achieving mean absolute errors and mean squared errors that are more than $40\%$ lower. The second proposed model, the Hybrid Quantum Sequence-to-Sequence neural network, once trained, predicts photovoltaic power with $16\%$ lower mean absolute error for arbitrary time intervals without the need for prior meteorological data, highlighting its versatility. Moreover, our hybrid models perform better even when trained on limited datasets, underlining their potential utility in data-scarce scenarios. These findings represent progress towards resolving time series prediction challenges in energy forecasting through hybrid quantum models, showcasing the transformative potential of quantum machine learning in catalyzing the renewable energy transition.
quantum machine learning, hybrid quantum neural network, quantum depth infused layer, quantum LSTM, solar energy, photovoltaic power, time series
I Introduction
Electricity generation prediction, especially for photovoltaic (PV) systems, is a crucial tool for renewable energy adoption sabri2022novel ; li2020review . The global economy must radically reduce emissions to stay within the $1.5\degree$C pathway (Paris Agreement) and the transition to renewable energy sources is necessary to achieve these objectives ellabban2014renewable . According to the International Energy Agency, solar PV’s installed power capacity is poised to surpass that of coal by 2027, becoming the largest in the world.
Accurate PV power forecasts are vital for multiple facets of the energy industry such as long-term investment planning, regulatory compliance for avoiding penalties, and renewable energy management across storage, transmission, and distribution activities. Several studies show that an increase in forecasting accuracy reduces electricity generation from conventional sources. Increased accuracy also reduces operating costs of systems through reducing the uncertainty of PV power generation anido2016value . They support improving the stability and sustainability of the power grid through optimizing power flow and counteracting solar power’s intermittent nature lee2021pv . Such predictions are foundational in increasing the economic viability and improving the adoption of solar energy as they inform pricing and economic dispatch strategies, bolster competitiveness and over time reduce reliance on reserve power. Additionally, they assist in managing energy storage effectively and integrating PV systems into the power grid bozorgavari2019two , which is essential for the enduring success of renewable energy solutions diagne2013review .
Traditional methods for predicting PV power have primarily relied on statistical models, machine learning algorithms, or a blend of both ahmed2019review . These approaches encompass a diverse toolkit, ranging from time series forecasting and artificial neural networks sabri2022novel ; gomez2020photovoltaic ; kaloop2021novel , to support vector machines kulkarni2018output ; shi2012forecasting , k-nearest neighbor methods rocha2022computational , and random forest models inman2013solar . However, the intermittent and non-linear nature of solar power generation, influenced by a wide range of meteorological factors, poses a significant challenge to the performance of these conventional models blaga2019current .
In light of these challenges, quantum machine learning (QML) emerges as a promising avenue. This rapidly evolving field, which melds the principles of quantum mechanics with classical machine learning Schuld_2014 ; biamonte2017quantum ; Dunjko.Briegel.2018 ; qml_review_2023 , can offer enhanced capabilities for improving the forecasting accuracy of time series tasks emmanoulopoulos2022quantum , including PV power generation sushmit2023forecasting . QML’s potential arises from quantum features like superposition and entanglement, promising exponential speedups in certain tasks schuld2018implementing . Moreover, QML algorithms produce inherently probabilistic results, aptly suited for prediction tasks, and they may potentially function within an exponentially larger search space, amplifying their efficacy Lloyd2013 ; 365700 ; Lloyd1996UniversalQS ; schuld2020circuit . Nonetheless, implementing quantum algorithms bears its own set of challenges, such as the need for error correction and sensitivity to external interference Aaronson2015 . Yet, in spite of these challenges, hybrid quantum-classical models, especially hybrid quantum neural networks (HQNNs), have showcased their potential in diverse industrial realms, including healthcare jain2022hybrid ; pharma ; steatosis_sagingalieva , energy kurkin2023forecasting ; sushmit2023forecasting , aerospace rainjonneau2023quantum , logistics haboury2023supervised and automotive sagingalieva2023hyperparameter industries.
Figure 1: The input to a hybrid quantum model is presented as a chronological data table, documenting hourly meteorological parameters including ambient temperature ($T_{a}$), module temperature ($T_{m}$), and solar irradiance ($I_{3}$, $I_{15}$), alongside the mean PV power output ($P$). The model is designed to leverage this data to generate predictions of PV power output for a short-term forecast aimed at near-future output, typically the next hour, and a long-term forecast that extends to a broader temporal horizon.
In this article, we present three types of hybrid quantum models as potential solutions for PV power forecasting. We assess the performance of our proposed models using a publicly accessible dataset, encompassing a comprehensive array of meteorological variables as well as hourly mean PV power measurements spanning a $21$-month period. This dataset, along with the data preprocessing and analytical methods employed, is described in detail in Section II.1.
Our first proposed HQNN architecture, articulated in Section II.2, incorporates classical fully connected layers with a vanilla variational repetitive quantum layer (VVRQ). Our second model, delineated in Section II.3, constitutes a hybrid quantum adaptation of the classical recurrent neural network, termed the Hybrid Quantum Long Short-Term Memory with quantum depth-infused layer (HQLSTM). While the first two models can predict the power for a certain hour ahead, the third model, presented in Section II.4, a Hybrid Quantum Sequence-to-Sequence Neural Network with quantum depth-infused layer, HQSeq2Seq, after training, is capable of forecasting PV power for arbitrary time intervals without requiring prior meteorological data.
Remarkably, despite having fewer parameters, our hybrid quantum models outperform their classical counterparts in terms of more accurate predictions, including when trained on a reduced dataset. We summarize our conclusions and outline future research directions in Section III.
II Results
The application of HQNNs in addressing time series prediction challenges, specifically in forecasting PV power output, offers several advantages. Primarily, their ability to operate within an exponentially larger computational search space enables them to efficiently capture intricate data patterns and relationships kordzanganeh2023exponentially . This feature not only enhances forecast accuracy biamonte2017quantum ; senokosov2024quantum but also streamlines the learning process, requiring fewer iterations for model optimization asel1 . Furthermore, the inherent capacity of quantum technologies to manage the uncertainty and noise ubiquitous in data offers more resilient and trustworthy predictions schuld2018implementing . This is particularly pertinent to power forecasting, given the inherent noise in meteorological data. Recent research also suggests that quantum models can be represented as partial Fourier series, positioning them as potential universal function approximators schuld2021effect , thereby broadening their applicability and efficacy in predictive tasks.
In terms of architecture, an HQNN is an amalgamation of classical and quantum components. The classical segments may consist of fully connected layers, convolutional layers, or recurrent layers, while the quantum segments are typically represented by variational quantum circuits (VQCs) or their contemporary modifications kordzanganeh2023parallel ; sedykh2024hybrid .
Figure 2: (a) Mean and standard deviation of the PV power value for each hour of the day. (b) Mean and standard deviation of the PV power value for each month of the year. The plot shows the PV power reaching maximal values in June and July. (c) Correlation matrix of input features. (d) Joint distribution of features.
II.1 Dataset
To investigate the advantages of hybrid quantum models, we selected a publicly accessible dataset malvoni2016data from a conventional generation plant situated in the Mediterranean region. This dataset not only provides comprehensive data but also enables benchmarking with results from various algorithms available in literature. A comparative analysis of our model’s predictions and those from the study by kaloop2021novel is provided in Section II.5. The dataset, presented as a numerical table showcased in Fig. 1, includes variables like: hourly mean ambient temperature, $T_{a}$; hourly mean module temperature, $T_{m}$; hourly mean solar irradiance recorded on two tilted planes with tilt angles of 3 and 15 degrees, $I_{3}$ and $I_{15}$; and hourly mean PV power, $P$, spanning $21$ months, accounting for more than $500$ days.
Beyond the scope of constructing models for predicting the output of PV panels, this dataset’s utility extends to other applications. It aids in planning distributed battery energy storage systems aghaei2019flexibility , devising novel energy collection systems kulkarni2018performance , and researching the degradation patterns of photovoltaic panels malvoni2017study . The dataset’s multifaceted applicability emphasizes its significance.
To ensure the validity and precision of the data, careful preprocessing was performed. We discovered approximately 20 anomalies in the original dataset. To maintain a continuous timeline, missing data points were replaced with the arithmetic mean of the preceding and succeeding day’s values. Additionally, data corresponding to the date “12/31/13” was excluded since it contained all-zero values, suggesting an error in the data collection. As a result, we obtained an uninterrupted dataset ranging from 4:00 AM on “3/5/12” to 12:00 AM on “12/30/13.”
Additional in-depth analysis of the dataset was also conducted for a more nuanced understanding. Fig. 2(a) delineates the hourly distribution of PV power across the entire period. As expected, peak PV power values occur during midday, while night time values plummet to zero. Fig. 2(b) portrays monthly PV power fluctuations, which are more volatile compared to daily patterns, likely attributable to the limited number of full-year periods in the dataset. Fig. 2(c) presents a correlation matrix for the dataset features, identifying solar irradiances $I_{3}$ and $I_{15}$ as the features most correlated with PV power. Finally, the joint distribution of dataset features depicted in Fig. 2(c) further confirms that solar intensity is the feature most highly correlated with PV power.
Figure 3: The architectures of: (a) Hybrid Quantum Neural Network with a VVRQ layer, (b) Hybrid Quantum Long Short-Term memory, (c) QDI layer used in HQLSTM model, (d) Hybrid Quantum Seq2Seq with QDI layer.
II.2 HQNN
This section introduces our first proposed model, referred to as the HQNN. As illustrated in Fig. 1, the model accepts weather data spanning $24$ consecutive hours as its input. The output is a prediction of the PV power for the upcoming $25$th hour. The HQNN presented in Fig. 3(a) is a combination of classical fully-connected layers, in our case with $120$, $17$ and $8$ neurons, and a VVRQ layer, which is a VQC, consisting of $q$ qubits and $d$ repetitions of variational layers, each distinguished by unique weights. The choice of $120$ neurons is methodical: the model ingests $5$ distinct features for each of the $24$ hours, resulting in a total of $120=5\times 24$. The determination of the remaining parameters stemmed from an extensive hyperparameter optimization process, detailed in the subsequent sections.
Initially, every qubit in the VVRQ layer is set to the state $\left|0\right\rangle$. We subsequently encode the classical data by converting it into rotation angles around one of the $X$, $Y$, $Z$ axes using $R_{x}$, $R_{y}$, $R_{z}$ gates, respectively. This conversion employs the angle embedding technique PhysRev.70.460 . For each qubit, the rotation angle, denoted by $x_{j}$, is determined by the $j$-th component of the input vector.
Following this, the variational layer is applied, which can either generate “basic” or “strong” entanglement. For the “basic” entanglement, each qubit undergoes a rotation by an angle $w_{j}^{i}$ around the $X$ axis, followed by a layer of CNOT gates Barenco_1995 . For the “strong” entanglement, each qubit is sequentially rotated by the angles $w^{(Z_{1})}_{ji}$, $w^{(Y_{2})}_{ji}$ and $w^{(Z_{3})}_{ji}$ around the $Z$, $Y$ and $Z$ axes, respectively. This sequence is then followed by a layer of CNOT gates. In both cases, the variables $i$ and $j$ play crucial roles in determining the operations. The variable $i$ signifies the particular wire to which the operation is applied, and it takes values from the set ${1,2,\dots,q}$. Meanwhile, the variable $j$ represents the number of variational layers and ranges from ${1,2,\dots,d}$.
Lastly, all the qubits are measured in Pauli-$Z$ basis, yielding the classical vector $v\in\mathbb{R}^{q}$. This output serves as input for a subsequent classical fully-connected layer. This layer processes information from $q$ neurons into $1$ neuron that predicts the power value.
The proposed HQNN model will be compared with its classical analog – a Multilayer Perceptron (MLP) that consists of $4$ fully connected layers with $120$, $32$, $3$, $3$, and $1$ neurons. The number of neurons in each layer was selected by a hyperparameter optimization procedure, detailed in the subsequent sections.
II.3 HQLSTM
This section presents a description of our second hybrid model – HQLSTM, which is a hybrid analog of the classical LSTM model hochreiter1997long , with which predictions will be compared in the following sections. LSTM architectures have garnered significant attention in the realm of time series forecasting, including in predicting PV power gao2023short ; sharadga2020time ; tovar2020pv . HQLSTM models have performed well on tasks such as natural language processing wang2022adverse , detecting software vulnerabilities akter2022automated , and predicting solar radiation yu2023prediction .
In this proposed model, we added a quantum layer to each of the LSTM gates chen2020quantum . Our implementation is depicted in Fig 3(b). The input to the model is:
  1. 1.
    The current step information, represented by a green circle, $x(t)$. This is a tensor of size $5$, reflecting the five features for an hour, which include meteorological data and the PV power itself.
  2. 2.
    The information from the previous step, denoted by a purple circle, $h(t-1)$. It consists of a tensor of size $h_{dim}$. For the initial step, this is simply a zero vector.
These inputs are processed through classical fully-connected layers to yield vectors with a uniform dimension of $4n_{q}$. These vectors are then concatenated through a bitwise addition operation.
Subsequently, this concatenated vector is partitioned into four distinct groups for the four gates of the LSTM cell. As illustrated in Fig. 3(b-c), each group is directed to the input of its corresponding quantum layer, symbolized by the Quantum Depth-Infused (QDI) square.
The outputs from QDI layers are transformed via classical fully-connected layers to standardize their dimensions to $h_{dim}$. Following this, similar to the classical LSTM, activation functions together with transformations appropriate to each of the four gates, are applied to the outputs originating from the quantum layers. This processing produces the new cell state and hidden state vectors, $C(t)$ and $h(t)$, respectively.
The process operates in a cyclical manner. For each iteration, the vector from the current time step, $x(t)$, and the hidden vector from the previous step, $h(t-1)$, serve as inputs to the HQLSTM. This iterative process is executed as many times as the input width; in our case, the input width equals $24$. Subsequently, all the hidden vectors are concatenated to produce a single composite vector. This vector is then processed through a fully-connected layer consisting of a single neuron, which outputs a value that predicts the PV power.
In our first proposed architecture, the HQNN, the quantum layer functioned as a vanilla layer, where variational layers were sequentially placed after the encoding layer. In contrast, in the HQLSTM approach, we used a QDI layer pharma as depicted in Fig. 3(c).
In order to increase the expressivity, variational layers are positioned both before and after the encoding layer. The first variational layer is illustrated with green rectangles, the encoding layer with blue rectangles, and the final variational layer with purple rectangles.
II.4 HQSeq2Seq
Here we present a hybrid version of the Sequence-to-Sequence (Seq2Seq) model, first introduced in sutskever2014sequence . Seq2seq models are widely used in natural language processing tasks kalchbrenner2013recurrent , where the length of input and output sequence is not predetermined and can be variable. We can also apply the principle of Seq2Seq models to the power prediction task mu2023improved . This means we can feed the neural network with time series of arbitrary length and prompt it to give us the forecast for any number of hours ahead. In this problem setting, the longer the input time series is, the better the model prediction is. Similarly, the shorter the required output length, the easier it is for the model to generate the forecast.
The seq2Seq model is a type of encoder-decoder model. The encoder is given the entire input sequence, which it uses to generate a context vector. This vector is used as an input hidden state for the decoder, so it literally provides it with “context,” according to which the decoder will generate the forecast. This means that the hidden dimensions of the encoder and the decoder must match.
The decoder creates the output sequence step by step. It starts with only the most recent entry. Based on this entry and the context vector, the decoder generates the second entry and appends it to the existing one. The obtained two-entry sequence is once again fed into the decoder to generate the third entry. The cycle then repeats until the length of the generated sequence matches the length requested by the user.
We create and compare two models with Seq2Seq architecture: the classical Seq2Seq and the hybrid model called HQSeq2Seq. Both of these models have identical LSTMs acting as encoders and decoders. In the classical model, the decoder’s hidden output vector is mapped to the “Power” value with a single linear layer, while in HQSeq2Seq it is processed by a QDI layer pharma .
In the QDI layer, instead of attempting to use a qubit for each feature asel1 , we employed the data re-uploading technique schuld2021effect ; data_reuploading . Specifically, we work with four qubits and structure them into a lattice of depth $4$ (depicted as a blue big rectangle in Fig. 3(d)). Each of our $16$ input features leading to the quantum layer is intricately encoded within this lattice. The first four features are mapped onto the initial depth, followed by the subsequent features in blocks of four. Encoding these classical features into the quantum domain, we perform “angle embedding” using $R_{z}$ gates. This operation effectively translates the input vector into a quantum state that symbolizes the preceding classical layer’s data.
Entangling variational layers, signified by purple squares, are interspersed between every encoding layer, ensuring optimal Fourier accessibility. Each variational layer has two components: rotations governed by trainable parameters, and sequential CNOT gates. The rotations are implemented by quantum gates that manipulate the encoded input in line with the variational parameters, while the CNOT gates entangle the qubits.
Each large blue bounding box encompasses a variational layer (small purple squares). Prior to all the encoding layers, we introduce a single variational layer (small green squares) for enhanced model representation. Consequently, the total weight count in the quantum segment of our network is $20$. In the measurement phase, all the qubits execute a CNOT operation where the first qubit is the target. This ensures that when the first qubit is measured in the $Y$-basis, the measurement outcome depends on the all qubits involved in the circuit. In this way, the output of the quantum layer serves as the power value prediction for a specific hour.
The input size of the encoder and decoder can differ, which is a substantial benefit. For instance, we can use all of the $5$ features to create a context vector, but request to generate the forecast for only $1$ feature. Exploiting this advantage, we will feed the Seq2Seq model with a window of all known features and only demand the power forecast. For the sake of simplicity, we will train both models with a fixed input and output length of $96$ hours and then vary the length in the testing stage.
Figure 4: Results of training and testing the HQNN, MLP, HQLSTM and LSTM models. (a-b) Training and testing history of the models shown by the dotted and solid lines, respectively. The filled space shows the standard deviation of that models averaged over the different testing subsets. (c) Histogram showing MAE and RMSE loss on the test dataset averaged over the results of 5 models trained on different subsets of training and test data. (d) The histogram shows MSE loss on the test dataset for MLP, HQNN, LSTM and HQLSTM models for the reduced training dataset size. (e) The histogram shows the RMSE loss on the test dataset for the Seq2Seq and HQSeq2Seq models for the reduced training dataset size. (f-g) The train and test learning curves for the Seq2Seq and HQSeq2Seq models. (h) An example of the classical and hybrid Seq2Seq models inference on the testing data. The models get $124$ hours of data as an input (before the dashed line) and give a forecast of the “Power” up to $137$ hours ahead (after the dashed line). The solid black line represents the ground truth value of the “Power” feature. The MSE/MAE errors in this particular example are $0.0052/0.0349$ for Seq2Seq and $0.0040/0.0292$ for HQSeq2Seq.
II.5 Training and results
In the study, six distinct models were employed for PV power prediction based on weather features: HQNN, MLP, HQLSTM, LSTM, HQSeq2Seq, and Seq2Seq. To train the models, the mean square error (MSE) was chosen as the loss function:
$\text{MSE}=\frac{1}{N}\sum\limits_{n=1}^{N}\left(x_{n}-y_{n}\right)^{2},$
where $N$ is number of predictions, $\vec{x}=\left(x_{1},x_{2},\dots,x_{N}\right)$ is the predicted PV power, and $\vec{y}=\left(y_{1},y_{2},\dots,y_{N}\right)$ is the actual PV power value.
To test the models, in addition to the MSE loss metric, we also used the mean absolute error (MAE), root mean squared error (RMSE) and variance account factor (VAF).
All the machine learning simulations for this study were conducted on CPUs, on the QMware cloud qmw_qmw_2022 ; kordzanganeh2023benchmarking device. The classical part of our modeling was structured using the PyTorch library PyTorch , while the quantum part was implemented using the PennyLane framework. Notably, PennyLane provides an assortment of qubit devices. For our requirements, we selected the lightning.qubit device which is a custom backend for simulating quantum state-vector evolution. To compute the gradients of the loss function relative to each parameter: for the classical components of our hybrid models, we employed the widely-recognized backpropagation algorithm backprop ; and for the quantum part, we used the adjoint method as highlighted in Refs. Luo2020yaojlextensible ; efficient_calculation .
Hyperparameters Range Best value
HQNN
number of neurons in
the second layer
$8-128$ 17
number of qubits $2-10$ 8
number of variational layers $1-10$ 7
embedding $R_{x}$, $R_{y}$, $R_{z}$ $R_{x}$
measurement $X$, $Y$, $Z$ $Z$
variational part basic, strongly basic
initial learning rate $1-1000\times 10^{-4}$ $3\times 10^{-2}$
MLP
number of neurons in
the first layer
$8-128$ 32
number of neurons in
the second layer
$8-128$ 3
number of neurons in
the third layer
$8-128$ 3
initial learning rate $1-1000\times 10^{-4}$ $1\times 10^{-2}$
HQLSTM
number of neurons in
the second layer
$3-25$ 20
dropout range $0-99$ 0.239
number of qubits $2-5$ 4
number of variational layers $1-5$ 1
number of quantum layers $1-5$ 3
initial learning rate $1-100\times 10^{-3}$ $0.52\times 10^{-2}$
LSTM
number of neurons in
the second layer
$3-25$ 21
dropout range $0-99$ 0.158
initial learning rate $1-100\times 10^{-3}$ $0.5\times 10^{-2}$
Table 1: The table summarizes the hyperparameter optimization process. It shows which hyperparameters are being optimized, the range over which they could be varied, and the best values found.
Model HQNN MLP HQLSTM LSTM
train, MAE $0.0458$ $0.0651$ $0.0382$ $0.0454$
test, MAE $0.0659$ $0.0887$ $0.0343$ $0.0570$
train, MSE $0.0059$ $0.0144$ $0.0056$ $0.0054$
test, MSE $0.0120$ $0.0204$ $0.0058$ $0.0096$
test, RMSE $0.1097$ $0.1428$ $0.0743$ $0.0937$
Table 2: Summary of the results for the proposed models. In a direct comparison, the HQNN outperforms the MLP in both training and testing losses, as evidenced across three critical metrics: RMSE, MAE, and MSE. Notably, the HQNN’s MSE is $41\%$ lower than that of the MLP. This demonstrates that the HQNN has more reliable forecasts, despite the fact that it has $1.8$ times fewer parameters. On the test dataset, the HQLSTM is $40\%$ better compared to the LTSM on the MAE and MSE metrics. It is also $21\%$ better on the RMSE metric, even though it has less than half as many weights. Overall, the HQLSTM emerges as the most precise model on all metrics. Specifically, it is $52\%$ more precise than the HQNN while having two times as few trainable parameters.
II.5.1 HQNN & MLP and HQLSTM & LSTM
Both the HQNN and its classical analog, the MLP, were trained for $20$ epochs. In contrast, the HQLSTM and its classical counterpart, the LSTM, were trained for over $50$ epochs. The Adam optimiser kingma2017adam from the PyTorch framework was used to update the parameters of the models in order to minimize their loss functions. The comprehensive training process, accompanied by the results, is shown in Fig. 4 and in Table 2.
In this study, we employed cross-validation as a fundamental technique to assess the performance of our models across distinct testing subsets. The application of cross-validation is pivotal to safeguard against potential data leakage from the training dataset into the testing dataset. To achieve this, a rigorous approach was adopted wherein a 24-hour time window, from each side of the subsets, was systematically excluded from the dataset.
Furthermore, we performed cross-validation by partitioning the dataset into training and testing sets in a $4:1$ ratio. This strategy was implemented to promote a comprehensive evaluation of our models, as we carried out model training and assessment on five distinct data splits. Subsequent averaging was used to consolidate the results obtained from these splits, and the mean values served as the primary metric for inter-model comparisons.
In a head-to-head comparison between the HQNN and the MLP, the former exhibits superior performance regarding training and testing losses across three key metrics: RMSE, MAE, and MSE. Specifically, the HQNN surpasses the MLP’s power prediction accuracy by $41\%$ as estimated by the MSE loss function, and by $26\%$ on the MAE loss, all while boasting $1.8$ times fewer parameters (2266& 3987226639872266\ \&\ 39872266 & 3987).
When comparing the HQLSTM and LSTM models, the former outperforms the latter in both training and testing losses across all three of the aforementioned metrics. Remarkably, the HQLSTM has a $40\%$ better predictive ability than the LSTM as assessed by the MSE and MAE metrics. It achieves this with less than half the number of parameters (1109& 2857110928571109\ \&\ 28571109 & 2857). Moreover, HQLSTMs are more resistant to overfitting, while classical LSTMs suffer from it.
In a broader comparison encompassing all four models, the HQLSTM emerges as the most precise model on all the metrics. Notably, it is $52\%$ more precise than the HQNN, while having half as many trainable parameters.
To achieve these results, we performed hyperparameter optimization using the Optuna optimizer optuna_link . The set of optimized parameters, the limits on their variation, and the best sets of hyperparameters for all our four models are presented in Table 1.
Moreover, we searched for external articles that refer to this dataset and found only one article that solves the $1$ hour ahead PV power prediction using neural networks kaloop2021novel . In comparison, our HQNN model is better than the model from this external article according to the VAF metric by $40\%$ (91& 65916591\,\&\,6591 & 65).
Further, to confirm that hybrid models train better including on a smaller dataset, an additional experiment was conducted where the volume of training data was intentionally reduced. The results are shown in Figure 4. The hybrid models performed better with less data, having lower losses and better prediction capabilities than the classical models.
II.5.2 HQSeq2Seq & Seq2Seq
After the preprocessing steps described in the Section II.1, the dataset spanned $12775$ hours from 3/5/12 4:55 AM to 8/19/13 10:00 AM for training, and $3194$ hours from 8/19/13 11:00 PM to 12/30/13 00:00 AM for testing.
Although the models are capable of being trained on sequences of arbitrary lengths, we chose to use sequences with a fixed $96$ hour length for simplicity. In this case, the encoder gets $96$ hours of all available features, while the decoder is asked to extrapolate only the “Power” feature of the data $96$ hours ahead. Training for $15$ epochs with the Adam optimizer (learning rate $0.001$) proved to be enough for the models to converge (Fig. 4 (f-g)).
As an example of inference, we pass the time series of the length different from $96$ into both models and prompt them to give us a forecast for $137$ hours ahead (Fig. 4 (h)). We can conclude that both models transition from the fixed sequence length to an arbitrary one quite well. It may even be possible to improve these results by introducing variable-length sequences into the training stage.
We also investigated the dependence of the test loss on the size of the training dataset for Seq2Seq and HQSeq2Seq, as shown in Fig. 4(e). As one can see, the test RMSE loss of the hybrid model is less for any size of training data, showing that the hybrid model has an advantage over the classical model, including on a substantially trimmed dataset.
III Discussion
In this work, we introduced three hybrid quantum approaches to time series prediction. The first two models allow one to predict the power of solar panels $1$ hour ahead using weather features from the previous $24$ hours. The third model is the most versatile since it enables longer-term, user-defined forecasts to be made.
The first approach is the HQNN, a combination of classical fully connected layers and a VVRQ layer, which is a quantum analog of such layers. We compared this hybrid model with its classical counterpart, the MLP, and demonstrated that, even though the HQNN has $1.8$ times fewer variational parameters, its predictive ability is $41\%$ better, as estimated by the MSE metric.
The second approach is the HQLSTM, a hybrid quantum analogue of the classical LSTM. Here, a QDI layer is inserted into each gate of the LSTM cell. This approach provides a $40\%$ improvement in prediction using the MAE and MSE metrics compared to its classical counterpart.
The HQLSTM was even an improvement on the HQNN despite having half as many weights.
The third approach is the hybrid Seq2Seq model, which consists of two LSTMs with a quantum layer at the end. This approach allows one to predict the PV power not only for an hour ahead, but for any number of hours ahead, without knowing the weather features in advance. The addition of the proposed QDI layer improves the accuracy of the predictions, reducing the MAE error by $16\%$ compared to a purely classical Seq2Seq model.
We conducted an additional experiment in which all our models were trained on a reduced dataset. It confirmed that hybrid models have better learning capabilities and lower losses than their classical counterparts when trained on any amount of data. This motivates the use of hybrid networks for applications where data collection is a complex task.
We also compared our models to a paper that solved the same problem using the same dataset. We found that our best HQLSTM is $40\%$ more accurate in predicting power using the VAF metric.
To fully unlock the potential of hybrid quantum approaches in time series prediction problems, further research and testing of models on other datasets is needed. More efficient optimization and training of VQCs, combined with higher fidelity and larger-scale quantum hardware, could lead to significantly greater performance improvements.
While this work was done on a public dataset with an emphasis on hybrid quantum models for better forecasting performance, the quality and source of data plays a crucial role in overall effectiveness in the real world, especially considering weather data. Accurate weather forecast is a crucial input into any high performing and useful PV prediction given its dynamism and influence on PV output. An interesting area of research is cloud prediction using satellite and weather data for geolocations directly impacting solar irradiance and therefore PV output. The added complexity could enhance the need for hybrid quantum models to increase computational efficiency and higher quality forecasts.
To summarize, our developments provide three hybrid quantum approaches for time series problems that demonstrate the possibility of combining classical and quantum methods. Our proposed models show improved performance compared to classical models with similar architectures when using fewer variational parameters. We believe that these results pave the way for further research in developing hybrid models that leverage the strengths of both classical and quantum computing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability All datasets used in this study are publicly available and are cited within the Dataset subsection of the paper. The machine learning models employed are fully described in the Results section, enabling independent reproduction of the results.